OpenStudy (anonymous):

Factor completely 2x2 - 6x – 36. 2(x - 3)(x + 6) 2(x - 6)(x + 3) (2x - 4)(x +9) (2x - 9)(x + 4)

OpenStudy (anonymous):

@iGreen @Lady.Liv1776 @Quan99 @_jazmineee @InExileWeTrust @january123

OpenStudy (anonymous):

I think it is A...

OpenStudy (anonymous):

Hmm yes, I believe you are right

OpenStudy (anonymous):

great thanks..do you mind helping me with one more?

OpenStudy (anonymous):

Sure! I'll try :)

OpenStudy (haseeb96):

the given equation is 2x2-6x-36=0 taking 2 as common so it would be as 2(x2-3x-18)=0 x2-3x-18=0 x2-6x+3x-18=0 x(x-6)+3(x-6)=0 (x-6)(x+3)

OpenStudy (haseeb96):

match your answer with it

OpenStudy (anonymous):

Determine the factors of x^2 - 6x - 8. Prime (x + 4)(x - 2) (x - 4)(x + 2) (x + 8)(x - 1)

OpenStudy (haseeb96):

is this equation correct?

OpenStudy (anonymous):

Meaning I typed the question right? Yes

OpenStudy (anonymous):

It is the second one

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

because 4 and 2 are factors of 8 that add to 6 and the middle term has to be negative

OpenStudy (haseeb96):

yes when we multiply them u got exact that equation which is given in the above question's equation.

OpenStudy (anonymous):

So it is (x + 4)(x - 2)

OpenStudy (anonymous):

No, it is (x-4) (x+2)

OpenStudy (anonymous):

oh....I thought you said B

OpenStudy (anonymous):

thank you though

OpenStudy (haseeb96):

x2 +2x -4x -8=0 x2 -2x-8=0 this is incorrect

OpenStudy (anonymous):

Oh, no, sorry! It's C :) haha

OpenStudy (haseeb96):

@jcbako u should again check it from your note book

OpenStudy (haseeb96):

C is also not the correct answer

OpenStudy (anonymous):

Im thinking Prime

OpenStudy (haseeb96):

did u check it?

OpenStudy (anonymous):

I tried all other answers and did not get the equation